OpenStudy (anonymous):

Write the equation of the line through (22, 4) that is parallel to the y=3 . Show your work.

OpenStudy (jamesj):

If you have two lines y = mx + c y = nx + d ...what must be true if the lines are parallel? I.e., what can you say about m, n, c or d?

OpenStudy (anonymous):

same slope

OpenStudy (jamesj):

hint: two parallel lines have the same slope

OpenStudy (jamesj):

Right, hence m and n is the same. With the line, y = 4, what is the slope of that line?

OpenStudy (anonymous):

the slope is 0 right

OpenStudy (jamesj):

So if we write y = 4 in the form y = mx + c, what is m, the slope?

OpenStudy (anonymous):

and its y=3

OpenStudy (jamesj):

Yes, sorry, y = 3 y = 3 is the same as writing y = 0x + 3 i.e., m = 0, c = 3. The slope is m = 0 So if we have a new line passing through (22,4), it must also have slope of zero. What line with slope zero passes through (22,4)?

OpenStudy (jamesj):

y = (something)x + (something)

OpenStudy (anonymous):

but if a slope is 0 there is no rise over run

OpenStudy (jamesj):

Yes, so it must be y = 0x + (something) i.e., y = something What must that something be so the line passes through (22,4)?

OpenStudy (anonymous):

it has to have the same slope?

OpenStudy (jamesj):

Yes, that is why the coefficient of x, the number before the x, is zero.

OpenStudy (anonymous):

but i dont get what the equation would be

OpenStudy (anonymous):

just y=0x+3?

OpenStudy (jamesj):

We know the slope must be zero. That is, for a straight line y = mx + c it must be that m = 0 That is, the line is of the form y = 0x + c i.e., y = 3 Yes, y = 0x + 3, is the first line. Now what then must be the equation of the line parallel to this that passes through (22,4)? It must have slope zero, hence we know the equation is of the form y = 0x + c for some value of c. What must that value of c be?

OpenStudy (anonymous):

ohhhhh its 4=0x+22

OpenStudy (anonymous):

or no?

OpenStudy (jamesj):

Nooo ... the equation you just wrote down implies that 4 = 22, which is clearly false. y = 0x + c and passes through (22,4), which means that (x,y) = (22,4) satisfies that equation.

OpenStudy (jamesj):

Substitute x = 22, y = 4 into the equation y = 0x + c and solve for c

OpenStudy (anonymous):

im so CONFUSED NOW :(

OpenStudy (anonymous):

OHHHHH OKAY!

OpenStudy (jamesj):



OpenStudy (anonymous):

C=4

OpenStudy (jamesj):

That dot is the point (22,4). What straight line passes through that point parallel to y = 4? Yes, y= 0x + 4 i.e., y = 4

OpenStudy (anonymous):

SO THAT WOULD BE MY EQQUATION RIGHT

OpenStudy (jamesj):

Yes, y = 4

OpenStudy (anonymous):

BUT IF ITS AN EQAUTION WHY CANT I RIGHT IT LIKE Y=0X+4

OpenStudy (jamesj):

You can write it as y = 0x + 4 if you like.

OpenStudy (jamesj):

But as 0x = 0 for all values of x, it is simpler to write y = 4

OpenStudy (anonymous):

oh i see! THANK YOOU SO MUCH! youre literally a better teacher than my own teacher

OpenStudy (jamesj):

thanks